# Supervised Graph Contrastive Learning for Gene Regulatory Network

### v3

@@ C.3 Extraction of Gene Label Data for BP/CC Tasks using Gene Ontology
In this study, for Biological Process (BP) classification and Cellular Component (CC) classification in downstream tasks, we performed multi-label annotation for each of the 975 genes constituting the GRN, based on terms obtained from the GO database. For BP labels, three categories whose importance is known were used: ’Metabolism’, ’Signal Transduction’, and ’Cellular Organization’ [31]. Similarly, for CC labels, four categories were used: ’Nucleus’, ’Mitochondrion’, ’Endoplasmic Reticulum’, and ’Plasma Membrane’ [32]. These categories were selected to cover the major functions of the GRN.
@@
    Batch retrieve terms associated with each gene using the MyGene.info API [33] and the OBO file (available at https://geneontology.org/docs/download-ontology/).
@@
    Aggregate multi-labels for the target genes using the GOOtools library [34].
@@
Table 5: Gene label distribution for high-level BP/CC categories (n = 975)
@@
C.3.1 Annotation by OncoKB
For cancer-related gene classification in downstream tasks, cancer-related genes were obtained from the OncoKB™ Cancer Gene List provided by OncoKB (Oncology Knowledge Base) [25]. Using the list of 1188 cancer-related genes provided as of April 30, 2025, positive labels were assigned to the 975 genes used in this study. As a result, 106 genes were labeled as cancer-related genes.
@@ D.1 Taxonomy of Gene Regulatory Network Estimation
Estimating accurate gene regulatory networks (GRNs) is crucial for elucidating cellular processes and disease mechanisms. Methods for computationally estimating GRNs from gene expression data can be categorized into correlation-based [35], mutual information-based [36], probabilistic graphical models-based such as Bayesian networks [37], and deep learning-based approaches [38]. The number of experimentally validated GRNs is limited. Therefore, GRN estimation methods need the ability to account for measurement errors and appropriately capture non-linear and multimodal interactions between genes while mitigating overfitting. Thus, we adopted a method combining Bayesian network estimation using multiple sampling and non-parametric regression [26] to identify patient- or sample-specific GRNs [8].
D.1.1 Details of Estimating GRNs
@@
Let $x_{1},x_{2},...,x_{n}$ be random variables for $n$ nodes, and let $\mathrm{pa}(i)$ be the set of parent nodes of the $i$-th node. In this case, a Bayesian network using B-spline curves [26] is defined as a probabilistic model decomposed into conditional distributions with parent nodes:
@@
The Bayesian network is estimated by learning the relationships with parent nodes based on the model described above and the parameters of the conditional distributions. In this study, the score function used for searching the structure of the Bayesian Network can be analytically derived using Laplace approximation [26]. Since the problem of finding a Directed Acyclic Graph (DAG) that maximizes this score is NP-hard, we performed structure search using a heuristic structure estimation algorithm, the greedy hill-climbing (HC) algorithm [39]. Furthermore, to ensure the reliability of the estimation results by the HC algorithm, we performed multiple sampling runs.
@@
Using the network and conditional probabilities learned here, we derive patient-specific GRNs [27]. The node set and edge set of the graph for a patient-specific GRN are defined by the network learned with gene expression levels as random variables $x_{1},\ldots,x_{n}$. Furthermore, the feature of the $i$-th node $X_{i}^{\mathcal{V}}$ is the gene expression level of each sample, and the feature of an edge from $i$ to $j$ (designated as the $k$-th edge) is the realization of the learned B-spline curve $X_{k}^{\mathcal{E}}=m_{ij}(x_{j})$. Patient-specific networks using such edge features have led to the discovery of subtypes that correlate more strongly with prognosis than existing subtypes [8]. Additionally, using differences in edge features between patients to extract patient-specific networks has been reported to contribute to the identification of novel diagnostic and therapeutic marker candidates in diseases such as idiopathic pulmonary fibrosis [40] and chronic nonbacterial osteomyelitis [41]
For GRN estimation, we used INGOR (version 0.19.0), a software that estimates Bayesian Networks based on B-spline regression, and executed it on the supercomputer Fugaku. INGOR is based on SiGN-BN [42], a software that similarly estimates Bayesian Networks using B-spline regression, and achieves faster estimation by optimizing parallel computation on Fugaku. In all GRN estimations, the number of sampling runs was set to 1000, and the threshold for adopting edges was set to 0.05. Since network estimation with a large amount of sample data can lead to Out of Memory errors, the upper limit of gene expression data used for network estimation was set to 3000 (especially for LINCS data). All other hyperparameters related to network estimation used the default settings of SiGN-BN. The number of Fugaku nodes and the required execution times are shown in Table 6. In addition, the number of parallel threads was set to 4 for estimations using TCGA data and 2 for LINCS data.
@@ Appendix E Experimental Setting
E.1 Hyperparameters Settings
[…]
@@ E.1 Hyperparameters Settings
E.2 Computational Environment and Computation Time
@@
E.3 Fine-tuning Settings
E.3.1 Graph-Level Task
@@
$h(t\mid x)=h_{0}(t)\exp\bigl(\beta^{\top}x\bigr{missing})$
Here, $h_{0}(t)$ is the baseline hazard, $x$ is the input variable, and $\beta$ is the regression coefficient to be learned. The prognosis estimation is performed by connecting this input variable $x$ to the graph NN and its head.
@@
E.3.2 Node-Level Task
@@
In Biological Process (BP) classification, three categories for each gene—"metabolism," "signal transduction," and "cellular organization"—are predicted as a multi-hot vector. In Cellular Component (CC) classification, four categories—"nucleus," "mitochondria," "endoplasmic reticulum," and "plasma membrane"—are predicted as a multi-hot vector. The model was structured using a sigmoid function for each category, and training was performed using binary cross-entropy for each respective category. Performance was evaluated using Subset Accuracy, Macro F1-score, and Jaccard Index as evaluation metrics. The Macro F1-score and Jaccard Index results are shown in Appendix F.
@@
In cancer-related gene classification, 106 genes defined as positive by OncoKB were labeled as "positive," and all other genes were labeled as "negative" for binary classification. A model was created to estimate negative and positive cases using a sigmoid function, and training was performed using binary cross-entropy. Performance was evaluated using Accuracy and F1-score as evaluation metrics. The F1-score results are shown in Appendix F.
@@ Appendix F Additional Results
As additional experimental results, we performed the following three analyses:
@@ F.1 Additional Evaluation Metrics
Tables 10 and 11 show the Macro F1-score and Jaccard index results for node-level tasks. Please note that for the cancer-related classification task, we evaluate only the F1-score because it involves binary data. Additionally, Table 12 shows the Macro F1-score for subtype classification in breast cancer. While our proposed method, SupGCL, did not individually achieve state-of-the-art results across all tasks and metrics, it demonstrated the most balanced performance overall.
@@ F.2 Additional Latent Space Analysis
Node-level embedding of other cancers:
Previously, in Result 2: Latent Space Analysis, we visualized the embedding space generated by pre-trained models on the breast cancer dataset (Figure 3). Analogous results for the lung cancer and colorectal cancer datasets are presented in Figure 4. Since subtype data were unavailable for the lung and colorectal cancer datasets, only node-level latent space visualizations are presented for these cancers. These results confirm that both GRACE and our proposed method, SupGCL, yield stable latent representations for these cancers, with no observed latent space collapse.
(a) Lung GRN Refer to caption (b) Colorectal GRN
Figure 4: t-SNE visualization of pre-trained embeddings on lung and colorectal cancer GRNs.
@@
To further investigate the characteristics of the node-level latent spaces presented in Result 2: Latent Space Analysis, we employed Principal Component Analysis (PCA). Figure 5 displays the PCA-projected latent spaces from pre-trained models on the breast cancer dataset, along with their corresponding explained variance ratios. For GraphCL, which previously exhibited tendencies towards latent space collapse, this analysis confirmed that its PCA explained variance ratio was overwhelmingly concentrated in the first principal component (PC1), accounting for 98.3%.

### v4

@@ C.3 Extraction of Gene Label Data for BP/CC Tasks using Gene Ontology
In this study, for Biological Process (BP) classification and Cellular Component (CC) classification in downstream tasks, we performed multi-label annotation for each of the 975 genes constituting the GRN, based on terms obtained from the GO database. For BP labels, three categories whose importance is known were used: ’Metabolism’, ’Signal Transduction’, and ’Cellular Organization’  (Paolacci et al. (2009)). Similarly, for CC labels, four categories were used: ’Nucleus’, ’Mitochondrion’, ’Endoplasmic Reticulum’, and ’Plasma Membrane’  (Costa et al. (2010)). These categories were selected to cover the major functions of the GRN.
@@
    Batch retrieve terms associated with each gene using the MyGene.info API (Xin et al. (2015)) and the OBO file (available at https://geneontology.org/docs/download-ontology/).
@@
    Aggregate multi-labels for the target genes using the GOATOOLS library  (Ashburner et al. (2000)).
@@
Table 6: Gene label distribution for high-level BP/$C$ categories (n = 975)
@@
C.4 Annotation by OncoKB
For cancer-related gene classification in downstream tasks, cancer-related genes were obtained from the OncoKB™ Cancer Gene List provided by OncoKB (Oncology Knowledge Base) (Chakravarty et al., 2017). Using the list of 1188 cancer-related genes provided as of April 30, 2025, positive labels were assigned to the 975 genes used in this study. As a result, 106 genes were labeled as cancer-related genes.
@@ D.1 Taxonomy of Gene Regulatory Network Estimation
Estimating accurate gene regulatory networks (GRNs) is crucial for elucidating cellular processes and disease mechanisms. Methods for computationally estimating GRNs from gene expression data can be categorized into correlation-based (Langfelder & Horvath (2008)), mutual information-based  (Margolin et al. (2006)), probabilistic graphical models-based such as Bayesian networks  (Friedman et al. (2000)), and deep learning-based approaches  (Shu et al. (2021)). The number of experimentally validated GRNs is limited. Therefore, GRN estimation methods need the ability to account for measurement errors and appropriately capture non-linear and multimodal interactions between genes while mitigating overfitting. Thus, we adopted a method combining Bayesian network estimation using multiple sampling and non-parametric regression (Imoto et al., 2002) to identify patient- or sample-specific GRNs (Nakazawa et al., 2021).
D.2 Details of Estimating GRNs
@@
Let $x_{1},x_{2},...,x_{n}$ be random variables for $n$ nodes, and let $\mathrm{pa}(i)$ be the set of parent nodes of the $i$-th node. In this case, a Bayesian network using B-spline curves (Imoto et al., 2002) is defined as a probabilistic model decomposed into conditional distributions with parent nodes:
@@
The Bayesian network is estimated by learning the relationships with parent nodes based on the model described above and the parameters of the conditional distributions. In this study, the score function used for searching the structure of the Bayesian Network can be analytically derived using Laplace approximation (Imoto et al., 2002). Since the problem of finding a Directed Acyclic Graph (DAG) that maximizes this score is NP-hard, we performed structure search using a heuristic structure estimation algorithm, the greedy hill-climbing (HC) algorithm  (Imoto et al. (2003)). Furthermore, to ensure the reliability of the estimation results by the HC algorithm, we performed multiple sampling runs.
@@
Using the network and conditional probabilities learned here, we derive patient-specific GRNs (Tanaka et al., 2020). The node set and edge set of the graph for a patient-specific GRN are defined by the network learned with gene expression levels as random variables $x_{1},\ldots,x_{n}$.
Furthermore, the feature of the $i$-th node $X_{i}^{\mathcal{V}}$ is the gene expression level of each sample, and the feature of an edge from $i$ to $j$ (designated as the $k$-th edge) is the realization of the learned B-spline curve $X_{k}^{\mathcal{E}}=m_{ij}(x_{j})$. Patient-specific networks using such edge features have led to the discovery of subtypes that correlate more strongly with prognosis than existing subtypes (Nakazawa et al., 2021). Additionally, using differences in edge features between patients to extract patient-specific networks has been reported to contribute to the identification of novel diagnostic and therapeutic marker candidates in diseases such as idiopathic pulmonary fibrosis  (Tomoto et al. (2024)) and chronic nonbacterial osteomyelitis  (Yahara et al. (2022))
For GRN estimation, we used INGOR (version 0.19.0), a software that estimates Bayesian Networks based on B-spline regression, and executed it on the supercomputer Fugaku. INGOR is based on SiGN-BN  (Tamada et al. (2011)), a software that similarly estimates Bayesian Networks using B-spline regression, and achieves faster estimation by optimizing parallel computation on Fugaku. In all GRN estimations, the number of sampling runs was set to 1000, and the threshold for adopting edges was set to 0.05. Since network estimation with a large amount of sample data can lead to Out of Memory errors, the upper limit of gene expression data used for network estimation was set to 3000 (especially for LINCS data). All other hyperparameters related to network estimation used the default settings of SiGN-BN. The number of Fugaku nodes and the required execution times are shown in Table 7. In addition, the number of parallel threads was set to 4 for estimations using TCGA data and 2 for LINCS data.
@@ Appendix E Experimental Setting
E.1 Validity of GNN Encoder
In this study, we consistently employed the Graph Transformer as the GNN encoder architecture. The choice of encoder is a critical factor that directly affects model performance, and its validity must be carefully assessed for our method. Our decision to adopt the Graph Transformer was based on prior work on GRNs (Liu et al., 2023) (Appendix E.1), where it achieved the best performance among various architectures. That study evaluated multiple GNN encoders—including GCN, GAT, TransformConv, SURGL, GPS, and GRACE—in the context of learning gene representations by integrating heterogeneous modalities such as scRNA-seq, scATAC-seq, and spatial transcriptomics. Their results showed that the Graph Transformer consistently outperformed other models, whereas GCN- and GAT-based models failed to sufficiently capture gene functional similarity across datasets. It should be noted, however, that this prior work focused on heterogeneous modalities, and is therefore not directly comparable to our setting, which relies exclusively on GRNs.
Motivated by these findings, we chose the Graph Transformer over GCN or GAT in our framework. Both GAT and Graph Transformer share the ability to incorporate real-valued edge features into node updates. However, in our experiments (using GRACE as an example), replacing the Graph Transformer with GAT sometimes led to degraded performance. The results are reported in Table 9.
Table 9: Comparison of downstream task performance with different encoders in GRACE (Breast Cancer GRN).
Model Setting Hazard (c-index) BP (Subset Accuracy)
GRACE (Graph Transformer, in paper) $0.642\pm 0.064$ $0.230\pm 0.051$
GRACE (GAT) $0.632\pm 0.038$ $0.241\pm 0.040$
E.2 Hyperparameters Settings
In this study, we used Optuna (Akiba et al., 2019), a Bayesian optimization tool, for hyperparameter search in pre-training. The search space for all models included the AdamW learning rate $\mathrm{lr}\in[10^{-5},10^{-3}]$, batch size $\mathrm{batch\_size}\in\{4,\,8\}$, and model-specific hyperparameters. Model-specific hyperparameters were the temperature parameter $\tau\in\{0.25,\,0.5,\,0.75,\,1.0\}$ for GraphCL, GRACE, and SupGCL, and the global-hop parameter $k\in\{1,2,3\}$ for SGRL. The graph embedding dimension was unified to 64 for all models. See Appendix F for a discussion on embedding dimension. For training the pre-trained models, the data was split into training and validation sets at an $8:2$ ratio, and the validation loss was used as the metric for various decisions. The optimal hyperparameters determined by actual hyperparameter tuning are shown in Table 10.
@@ E.2 Hyperparameters Settings
E.3 Computational Environment and Computation Time
@@
E.4 Fine-tuning Settings
E.4.1 Graph-Level Task
@@
[…]
Here, $h_{0}(t)$ is the baseline hazard, $\bm{x}$} is the input variable, and $\bm{\beta}$ is the regression coefficient to be learned. The prognosis estimation is performed by connecting this input variable $\bm{x}$} to the graph NN and its head.
@@
E.4.2 Node‐Level Task
@@
In Biological Process (BP) classification, three categories for each gene—”metabolism,” ”signal transduction,” and ”cellular organization”—are predicted as a multi-hot vector. In Cellular Component (CC) classification, four categories—”nucleus,” ”mitochondria,” ”endoplasmic reticulum,” and ”plasma membrane”—are predicted as a multi-hot vector. The model was structured using a sigmoid function for each category, and training was performed using binary cross-entropy for each respective category. Performance was evaluated using Subset Accuracy, Macro F1-score, and Jaccard Index as evaluation metrics. The Macro F1-score and Jaccard Index results are shown in Appendix F.
@@
In cancer-related gene classification, 106 genes defined as positive by OncoKB were labeled as ”positive,” and all other genes were labeled as ”negative” for binary classification. A model was created to estimate negative and positive cases using a sigmoid function, and training was performed using binary cross-entropy. Performance was evaluated using Accuracy and F1-score as evaluation metrics. The F1-score results are shown in Appendix F.
@@ Appendix F Additional Results
As additional experimental results, we performed the following six analyses:
@@
  4. 4.
    Statistical significance testing for the performance metrics across downstream tasks.
  5. 5.
    Comparison with Augmentation-Adjustment Methods and Augmentation-Free.
  6. 6.
    Evaluation on Link Prediction.
Furthermore, an analysis of the robustness of SupGCL against dataset changes is provided in Appendix H.
@@ F.1 Additional Evaluation Metrics
Tables 12 and 13 show the Macro F1-score and Jaccard index results for node-level tasks. Please note that for the cancer-related classification task, we evaluate only the F1-score because it involves binary data. Additionally, Table 14 shows the Macro F1-score for subtype classification in breast cancer. While our proposed method, SupGCL, did not individually achieve state-of-the-art results across all tasks and metrics, it demonstrated the most balanced performance overall.
@@ F.2 Additional Latent Space Analysis
Node-level latent Space:
Previously, in Result 3: Latent Space Analysis, we visualized the graph-level embedding space generated by pre-trained models(Figure 3). Node-level results for the breast, lung, and colorectal cancer datasets are presented in Figure 4. Since subtype data were unavailable for the lung and colorectal cancer datasets, only node-level latent space visualizations are presented for these cancers.
These results confirm that both GRACE and our proposed method, SupGCL, yield stable latent representations for these cancers, with no observed latent space collapse. A more detailed analysis on latent space collapse is provided in the next section. Furthermore, a more detailed analysis of the latent space from a biological perspective is provided in Appendix I
Figure 4: t-SNE visualization of pre-trained embeddings on breast, lung, and colorectal cancer GRNs.
@@
To further investigate the characteristics of the node-level latent spaces presented in Result 3: Latent Space Analysis, we employed Principal Component Analysis (PCA). Figure 5 displays the PCA-projected latent spaces from pre-trained models on the breast cancer dataset, along with their corresponding explained variance ratios. For GraphCL, which previously exhibited tendencies towards latent space collapse, this analysis confirmed that its PCA explained variance ratio was overwhelmingly concentrated in the first principal component (PC1), accounting for 98.3%.